### v1

@@ 8 Coupling
Simulating actual biological behavior involves complex coupling between different equations. Given we make heavy use of post-processing techniques, we couple the various solvers in a staggered way. We have then the option to choose between explicit and implicit staggering. Both approaches have shown to be effective in the literature [84, 65, 64, 51]. In [7, Appendix E], where fluid deformable membranes with phase separation are considered, both procedures show very similar results even in the most complex scenarios. Moreover, when implicit staggering is chosen, the authors of [7] claim that convergence is achieved in four steps on average. Strong of these results we proceed in coupling our solvers.
Remark 8.1.
An alternative option is to use a monolithic approach and couple the solvers in a unique system, avoiding staggering. In this way there is no need for sub-iterations and memory allocation is reduced. On the other side, the solution is usually only available for ad-hoc problems and not flexible. Together with this, eventual non-linearities have to be linearized. One can find an example of such a choice in [85], where wetting dynamics of liquid droplets on deformable membranes is simulated.
@@ 8.1 Numerical results for coupled mean curvature and ADR equations
We begin by testing the convergence properties of our coupled solver. To do so we pick the coupling example presented in [67]. The system which is solved is the following [67, p.686]:
(68a) $\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f(t,{\boldsymbol{x}}),$
(68b) ${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ $=(\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma},$
with parameters $\alpha,\beta,\delta\in\mathbb{R}^{+}$ and ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$. Convergence studies are perfomed choosing $f,g$ such that the exact solution for $u$ is $u(t,{\boldsymbol{x}})=x_{1}x_{2}e^{-6t}$. The geometry is chosen to be a sphere whose radius evolves following the law
(69) $R(t)=\frac{r_{0}r_{K}}{r_{K}e^{-kt}+r_{0}(1-e^{-kt})},$
@@
To simulate (68a) we can use what has been introduced in Section 5 and add a right-had side coupling. Since we chose $\alpha=0$, we can simulate (68b) by simplifying (65). The result is a weighted mean curvature solver that, in this very case, takes the form
(70a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-\beta({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}}$ $=((\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}},$
(70b) $({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ $=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}.$
We perform the test using implicit staggering. The results of the convergence studies are shown in Figure 11, where the same norm as in Equation 67 is used to study the results. One can see how the expected convergence is achieved in both space and time.
Figure 11: Convergence studies for mcadrSolver 4 and problem (68).
To qualitatively visualize the effect of the mesh redistribution and ADR stabilization, the example in [67, Sec. 11.2] is also reproduced. The simulation is derived from a proposed model for tumor growth. The coupled system is as follows:
(71a) $\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f_{1}(u,w),$
(71b) $\partial^{\bullet}w+w\nabla_{\Gamma}\cdot{\boldsymbol{v}}-D_{c}\Delta_{\Gamma}w$ $=f_{2}(u,w),$
(71c) ${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\nabla_{\Gamma}\mathbf{Id}_{\Gamma}$ $=\delta u{\boldsymbol{n}}_{\Gamma},$
@@
(72) $f_{1}(u,w)=\gamma(a-u+u^{2}w),\quad f_{2}(u,w)=\gamma(b-u^{2}w).$
@@
As discussed in [67], the nonlinear system composed by (71a) and (71b) is solved until $t=5$ without coupling it to the mesh evolution, i.e. ignoring (71c). After that, the full system is evolved until $t=8$. Implicit staggering is used. We thus have the following solvers:
@@
In Figure 12 we compare the mesh quality at the final time $t=8$. It is evident that the redistribution step keeps the mesh well-behaved while maintaining solution accuracy.
@@
Figure 12: Simulation at $t=8$ for the model (71) [67, Sec. 11.2] and the implicitly staggered schemes in List 8.1.
@@ 8.2 Numerical results for coupled Helfrich flow and Cahn-Hilliard euqations
We proceed now to test the coupling for more complicated systems and geometry evolution to explore the potential of our framework. The overdamped limit model for fluid deformable two-component membranes presented in [7, p.A41-33] is considered. The discrete scheme (65) is augmented with a Lagrange multiplier to maintain the inextensibility constraint $\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{v}}_{h}^{n}=0$. The modified version reads: Given ${\boldsymbol{y}}^{n}_{h},{\boldsymbol{\kappa}}^{n}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ find ${\boldsymbol{d}}^{n+1}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d}$, ${\boldsymbol{y}}^{n+1}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ and $\lambda_{h}^{n+1}\in V_{h}(\Gamma_{h}^{n})$ such that
(73a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}+(\lambda_{h}^{n+1}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})$
@@
(73b) $\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$
@@
(73c) $(\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}\mu_{h})_{\Gamma_{h}^{n}}+(|{\boldsymbol{\kappa}}_{h}^{n}|^{2}\lambda_{h}^{n+1}\,,\,\mu_{h})_{\Gamma_{h}^{n}}$
@@
(74) $\qquad+(({\boldsymbol{y}}^{n}_{h}\cdot{\boldsymbol{\kappa}}^{n}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{f}}_{h}^{n+1}\,,\,{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}.$
@@
    (75) ${\boldsymbol{f}}_{h}^{n+1}=\mathcal{E}_{CH}(u_{h}^{n+1}){\boldsymbol{\kappa}}_{h}^{n}-\sigma\epsilon\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1}\mathbb{H}_{h}^{n}\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1},$
    where ${\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is a certain vertex normal, see [11] for details, and $\mathbb{H}_{h}^{n}=-\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is the discrete extended Weingarten map.
@@
    (76) ${\boldsymbol{v}}_{h}^{\top}=-\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1},\quad{\boldsymbol{v}}_{h}^{\perp}={\boldsymbol{d}}_{h}^{n+1}/\tau.$
@@
(77) $u_{0}({\boldsymbol{x}})=\cos(\pi x_{1})\cos(\pi x_{2})\cos(\pi x_{3}).$
The simulation is run with variable elasticity modulus $\gamma_{W}\in\{0.5,~0.1,~0.02\}$ following the parameters in [7]. The timestep is varied accordingly with values $\tau\in\{10^{-4},~10^{-4},2.5\cdot 10^{-5}\}$ respectively. The simulation is stopped once the area difference between initial and evolved surface differs by more than $10\%$. Results are presented in Figure 13, where it can be seen that the characteristic bulging dynamics is correctly recovered.
@@
Figure 13: (a) Initial setup for problem the problem 77. (b)-(d) Mesh for different elasticity modulus. The simulation is shown at end time $t=1$ or at last timestep before inextensibility threshold is violated.

### v2

@@ 8 Coupling
Simulating actual biological behavior involves complex coupling between different equations. Given we make heavy use of post-processing techniques, we couple the various solvers in a staggered way. We have then the option to choose between explicit and implicit staggering. Both approaches have shown to be effective in the literature Mercker et al. [2013], Khanwale et al. [2023, 2020], Garcke et al. [2025]. In [Bachini et al., 2023, Appendix E], where fluid deformable membranes with phase separation are considered, both procedures show very similar results even in the most complex scenarios. Moreover, when implicit staggering is chosen, the authors of Bachini et al. [2023] claim that convergence is achieved in four steps on average. Strong of these results we proceed in coupling our solvers.
Remark 9.
An alternative option is to use a monolithic approach and couple the solvers in a unique system, avoiding staggering. In this way there is no need for sub-iterations and memory allocation is reduced. On the other side, the solution is usually only available for ad-hoc problems and not flexible. Together with this, eventual non-linearities have to be linearized. One can find an example of such a choice in Mokbel et al. [2024], where wetting dynamics of liquid droplets on deformable membranes is simulated.
@@ 8.1 Numerical results for coupled mean curvature and ADR equations
We begin by testing the convergence properties of our coupled solver. To do so we pick the coupling example presented in Barreira et al. [2011], Kovács et al. [2017]. The system which is solved is the following [Kovács et al., 2017, p.686]:
$\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f(t,{\boldsymbol{x}}),$ (69a)
${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ $=(\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma},$ (69b)
with parameters $\alpha,\beta,\delta\in\mathbb{R}^{+}$ and ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$. Convergence studies are performed choosing $f,g$ such that the exact solution for $u$ is $u(t,{\boldsymbol{x}})=x_{1}x_{2}e^{-6t}$. The geometry is chosen to be a sphere whose radius evolves following the law
$R(t)=\frac{r_{0}r_{K}}{r_{K}e^{-kt}+r_{0}(1-e^{-kt})},$ (70)
@@
To simulate (69a) we can use what has been introduced in Section 5 and add a right-hand side coupling. Since we chose $\alpha=0$, we can simulate (69b) by simplifying (66). The result is a weighted mean curvature solver that, in this very case, takes the form
$\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-\beta({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}}$ $=((\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}},$ (71a)
$({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ $=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}.$ (71b)
We perform the test using implicit staggering. The results of the convergence studies are shown in Figure 11, where the same norm as in (68) is used to study the results. One can see how the expected convergence is achieved in both space and time.
Figure 11: Convergence studies for mcadrSolver 4 and problem (69). As expected, first and second order convergence are achieved in time and space, respectively.
To qualitatively visualize the effect of the mesh redistribution and ADR stabilization, the example in [Kovács et al., 2017, Sec. 11.2] is also reproduced. The simulation is derived from a proposed model for tumor growth. The coupled system is as follows:
$\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f_{1}(u,w),$ (72a)
$\partial^{\bullet}w+w\nabla_{\Gamma}\cdot{\boldsymbol{v}}-D_{c}\Delta_{\Gamma}w$ $=f_{2}(u,w),$ (72b)
${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\nabla_{\Gamma}\mathbf{Id}_{\Gamma}$ $=\delta u{\boldsymbol{n}}_{\Gamma},$ (72c)
@@
$f_{1}(u,w)=\gamma(a-u+u^{2}w),\quad f_{2}(u,w)=\gamma(b-u^{2}w).$ (73)
@@
As discussed in Kovács et al. [2017], the nonlinear system composed by (72a) and (72b) is solved until $t=5$ without coupling it to the mesh evolution, i.e. ignoring (72c). After that, the full system is evolved until $t=8$. We study the following solvers:
@@
In Figure 12 we compare the mesh quality at the final time $t=8$. It is evident that the redistribution step keeps the mesh well-behaved while maintaining the solution accuracy.
@@
Figure 12: Simulation at $t=8$ for the model (72) [Kovács et al., 2017, Sec. 11.2] and the implicitly staggered schemes in List 8.1.
We can construct a bulk-surface problem by modifying the system in (72) with the following one
$\partial^{\bullet}u+u\nabla\cdot{\boldsymbol{v}}-\Delta u$ $=f_{1}(u,w)\quad\text{ in }\Omega,$ (74a)
$\partial^{\bullet}w+w\nabla\cdot{\boldsymbol{v}}-D_{c}\Delta w$ $=f_{2}(u,w)\quad\text{ in }\Omega,$ (74b)
${\boldsymbol{v}}-\alpha\Delta_{\partial\Omega}{\boldsymbol{v}}-\beta\nabla_{\partial\Omega}\mathbf{Id}_{\partial\Omega}$ $=\delta u{\boldsymbol{n}}_{\partial\Omega}\quad\text{ on }\partial\Omega.$ (74c)
The non-linear couplings and the parameters are kept the same as in (73) and Table 8.1. The initial geometry is a circle of unitary radius and the velocity prescribed by (74c) is extended in the bulk harmonically as described in Section 4. Results are presented in Figure 13, where it can be observed how the physics is correctly resolved while maintaining good mesh nodes distribution.
(a) mcadrSolver 1
(b) mcadrSolver 2
Figure 13: Simulation at $t=8$ for the model (74) and the implicitly staggered schemes in List 8.1.
@@ 8.2 Numerical results for coupled Helfrich flow and Cahn-Hilliard euqations
We proceed now to test the coupling for more complicated systems and geometry evolution to explore the potential of our framework. The overdamped limit model for fluid deformable two-component membranes presented in [Bachini et al., 2023, p.A41-33] is considered. The discrete scheme (66) is augmented with a Lagrange multiplier to maintain the inextensibility constraint $\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{v}}_{h}^{n}=0$. The modified version reads: Given ${\boldsymbol{y}}^{n}_{h},{\boldsymbol{\kappa}}^{n}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ find ${\boldsymbol{d}}^{n+1}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d}$, ${\boldsymbol{y}}^{n+1}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ and $\lambda_{h}^{n+1}\in V_{h}(\Gamma_{h}^{n})$ such that
$\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}+(\lambda_{h}^{n+1}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})$ (75a)
@@
$\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ (75b)
@@
$(\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}\mu_{h})_{\Gamma_{h}^{n}}+(|{\boldsymbol{\kappa}}_{h}^{n}|^{2}\lambda_{h}^{n+1}\,,\,\mu_{h})_{\Gamma_{h}^{n}}$ (75c)
@@
$\qquad+(({\boldsymbol{y}}^{n}_{h}\cdot{\boldsymbol{\kappa}}^{n}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{f}}_{h}^{n+1}\,,\,{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}.$ (76)
@@
    ${\boldsymbol{f}}_{h}^{n+1}=\mathcal{E}_{CH}(u_{h}^{n+1}){\boldsymbol{\kappa}}_{h}^{n}-\sigma\epsilon\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1}\mathbb{H}_{h}^{n}\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1},$ (77)
    where ${\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is a certain vertex normal, see Barrett et al. [2007] for details, and $\mathbb{H}_{h}^{n}=-\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is the discrete extended Weingarten map.
@@
    ${\boldsymbol{v}}_{h}^{\top}=-\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1},\quad{\boldsymbol{v}}_{h}^{\perp}={\boldsymbol{d}}_{h}^{n+1}/\tau.$ (78)
@@
$u_{0}({\boldsymbol{x}})=\cos(\pi x_{1})\cos(\pi x_{2})\cos(\pi x_{3}).$ (79)
The simulation is run with variable elasticity modulus $\gamma_{W}\in\{0.5,~0.1,~0.02\}$ following the parameters in Bachini et al. [2023]. The timestep is varied accordingly with values $\tau\in\{10^{-4},~10^{-4},2.5\cdot 10^{-5}\}$ respectively. The simulation is stopped once the area difference between initial and evolved surface differs by more than $10\%$. Results are presented in Figure 14, where it can be seen that the characteristic bulging dynamics is correctly recovered. As already noted in 7.2.1 and Figure 10, the pearlings’ mesh appears to be very coarse due to the extreme deformation. This is characteristic of the employed mesh redistribution algorithm presented in Section 4. As already noted in Garcke et al. [2025], alternative choices are possible that might lead to better results Sauer [2014], Hu and Li [2022], Porrmann et al. [2025], Sauer [2025], Zhang and Huang [2025], Duan [2025]. Our setting is flexible enough to allow to switch between various options. We repeat the test in 7.2.1 substituting the surface mesh redistribution in Section 4 Duan and Li [2024] with the minimal deformation rate (MDR) redistribution of Hu and Li [2022] presented in (34). The results are shown in Figure 15. It can be seen how using the MDR technique leads, in this very specific case, to better results with respect to Figure 14. This said, we found the performance of the different redistributions to be case-specific. For example, one can notice how, despite the enhanced node distribution, the inextensibility constaint for Figure 15(c) is violated at an earlier time with respect to Figure 14(c).
@@
Figure 14: (a) Initial setup for the problem 79. (b)-(d) Mesh for different elasticity modulus. The simulation is shown at end time $t=1$ or at last timestep before inextensibility threshold is violated.
(b) $\gamma_{W}=0.5,~t=1$
(c) $\gamma_{W}=0.1,~t=0.45$
(d) $\gamma_{W}=0.02,t=0.2375$
Figure 15: (a) Initial setup for the problem 79. (b)-(d) Mesh for different elasticity modulus. The simulation is shown at end time $t=1$ or at last timestep before inextensibility threshold is violated.